# Polynomial invariants for $3$-dimensional Leibniz algebras 111 The work is supported by the FCT 2023.08031.CEECIND, UIDB/00212/2020, UIDP/00212/2020, and FAPESP 2023/17918-2

### v1

@@ 3. $\bullet$
    the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$, $\mathcal{L}_{9}$ are not nilpotent for all $\lambda\in\mathbb{F}$.
@@
For the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$ the elements $(\cdots((e_{2}e_{3})e_{3})\cdots)e_{3}$ are non-zero. Similarly, $(\cdots((e_{1}e_{3})e_{3})\cdots)e_{3}\neq 0$ in $\mathcal{L}_{9}$. Since $(e_{3}e_{3})e_{3}=e_{2}\neq 0$ in $\mathcal{L}_{8}$, we can see that ${\rm ncl}(\mathcal{L}_{8})=4$. The cases of $\mathcal{L}_{4}^{\lambda}$ and $\mathcal{L}_{5}$ are trivial. ∎
@@
$\mathcal{L}\simeq\mathcal{L}_{1}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=2.$
$\mathcal{L}\simeq\mathcal{L}_{2}^{\lambda\neq 0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=(\lambda-1)\mathop{\rm tr}(\chi_{1}\chi_{0}).$
$\mathcal{L}\simeq\mathcal{L}_{2}^{0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=4.$
$\mathcal{L}\simeq\mathcal{L}_{3}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=-\mathop{\rm tr}(\chi_{1}\chi_{0}).$
$\mathcal{L}\simeq\mathcal{L}_{6}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=0.$
$\mathcal{L}\simeq\mathcal{L}_{7}^{\lambda\neq 0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=(1+2\lambda)\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$
$\mathcal{L}\simeq\mathcal{L}_{7}^{0}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=3.$
$\mathcal{L}\simeq\mathcal{L}_{9}$Longleftrightarrow $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$

### v2

@@ 3. $\bullet$
    the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{{\lambda\neq 0}}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$, $\mathcal{L}_{9}$, $\mathcal{L}_{10}$ are not nilpotent for all $\lambda\in\mathbb{F}$.
@@
For the algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{{\lambda\neq 0}}$, $\mathcal{L}_{3}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$ the elements $(\cdots((e_{2}e_{3})e_{3})\cdots)e_{3}$ are non-zero. Similarly, $(\cdots((e_{1}e_{3})e_{3})\cdots)e_{3}\neq 0$ in $\mathcal{L}_{9}$ and in $\mathcal{L}_{10}$. Since $(e_{3}e_{3})e_{3}=e_{2}\neq 0$ in $\mathcal{L}_{8}$, we can see that ${\rm ncl}(\mathcal{L}_{8})=4$. The cases of $\mathcal{L}_{4}^{\lambda}$, $\mathcal{L}_{5}$ and $\mathcal{L}_{11}$ are trivial. ∎
@@
$\mathcal{L}\simeq\mathcal{L}_{1}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=2.$
$\mathcal{L}\simeq\mathcal{L}_{2}^{\lambda\neq 0}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=(\lambda-1)\mathop{\rm tr}(\chi_{1}\chi_{0}).$
$\mathcal{L}\simeq\mathcal{L}_{3}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})\neq 0,$ $d=3,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=-\mathop{\rm tr}(\chi_{1}\chi_{0}).$
$\mathcal{L}\simeq\mathcal{L}_{6}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})=0.$
$\mathcal{L}\simeq\mathcal{L}_{7}^{\lambda\neq 0}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=(1+2\lambda)\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$
$\mathcal{L}\simeq\mathcal{L}_{7}^{0}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=3.$
$\mathcal{L}\simeq\mathcal{L}_{9}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=2,$ $\mathop{\rm tr}(\chi_{0}\chi_{1})\neq 0,\;\;\mathop{\rm tr}((\chi_{0}\chi_{1})\chi_{1})=\mathop{\rm tr}(\chi_{0}\chi_{1})^{2}.$
$\mathcal{L}\simeq\mathcal{L}_{10}$ $\Longleftrightarrow$ $\mathop{\rm tr}(\chi_{1}\chi_{0})=0,$ $d=4.$
Acknowledgement
We would like to express our sincere gratitude to the anonymous referee for the valuable and constructive comments, which have helped us improve the quality and clarity of our paper.